OpenStudy (nairatrebilus):

Let f(x) = x^3 + 2x^2 + 5x + 1, and let g be the inverse function of f. What is the value of g '(9)?

myininaya (myininaya):

I think I responded on one of your questions.

myininaya (myininaya):

I don't really want to type the same thing I did there here.

OpenStudy (aum):

\[ \text{If g(x) is the inverse fnction of f(x), then:}\\ g'(x)=\frac{1}{f'(g(x))} \\ g'(9)=\frac{1}{f'(g(9))} ~~~\text{------ (1)}\\ f(x) = x^3 + 2x^2 + 5x + 1 \\ \text{Normally we are given x, and we find y. }\\ \text{In inverse function we are given y and we need to find x.} \\ \text{To find g(9), set y = 9 and find x.} \\ 9 = x^3 + 2x^2 + 5x + 1 \\ x^3 + 2x^2 + 5x - 8 = 0 \\ \text{Rational roots theorem: Factors of 8 are: }\pm(1, 2, 4, 8). \\ \text{Plug each factor into }x^3 + 2x^2 + 5x - 8 = 0 \text{ and see which one}\\ \text{makes it zero. That x value will be g(9).} \\ \text{Then find f '(x) and plug it into equation (1) } \\ \]